In the Bohr model of hydrogen atom, the electron follows a circular orbit centred on


the nucleus containing a proton. The motion of the electron along the circular orbit


constitutes a current. Calculate the magnetic field produced by the orbiting electron


at the site of the proton.

1
Expert's answer
2022-02-28T10:41:43-0500

"v=e\sqrt{\frac k{m_1r}},"

"B=\frac{m_2v}{er}=\sqrt{\frac{km_2^2}{m_1r^3}}=8.4~T."